OpenStudy (anonymous):

Which of the following represents 4x to the 5/6 power in radical form?

OpenStudy (zzr0ck3r):

\[\sqrt[6]{(4x)^5}\]

OpenStudy (anonymous):

Thanks

OpenStudy (zzr0ck3r):

np

OpenStudy (skullpatrol):

Failure to follow the guidelines may lead to your suspension or banning from the site. If you see any violations, please report them by using a report abuse link. OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution.

OpenStudy (ash2326):

@zzr0ck3r Please explain how did you get this answer?

OpenStudy (zzr0ck3r):

or you could ? wow...

OpenStudy (zzr0ck3r):

\[a^{\frac{b}{c}}=\sqrt[c]{a^b}\]

OpenStudy (zzr0ck3r):

how is that any different than \[(4x)^{\frac{5}{6}}=\sqrt[6]{(4x)^5}\] @ash2326 ?

OpenStudy (zzr0ck3r):

its the exact same thing, but people hate a,b,c.....at this point in their math career.

OpenStudy (ash2326):

Yes, you're right. It's not rocket science. I just wanted you to explain a bit and not hand in the answer directly